OpenStudy (anonymous):

log3(a-2)+log3(a-2)=2

OpenStudy (anonymous):

it is log base of 3

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf log_3(a-2)+log_3(a-2)=2 }\) \(\LARGE\color{blue}{ \bf log_3(a-2)^2=2 }\) \(\LARGE\color{blue}{ \bf log_3(a-2)^2=2~log_33 }\) \(\LARGE\color{blue}{ \bf log_3(a-2)^2=log_33^2 }\)

OpenStudy (solomonzelman):

can you finish the problem ?

OpenStudy (anonymous):

where did you get 2 log33

OpenStudy (solomonzelman):

\[hint:~~~\log_aB=\log_aC~~~~~~there fore~~ B=C\] log base 3 of 3 is a magic one :)

OpenStudy (anonymous):

im not catching up, im sorry

OpenStudy (anonymous):

I think I got it! just remembered this! If y=b^x then log b y=x thanks!